### v1

# LEAD: Towards Learning-Based Equity-Aware Decarbonization in Ridesharing Platforms

Abstract
Ridesharing platforms such as Uber, Lyft, and DiDi have grown in popularity due to their on-demand availability, ease of use, and commute cost reductions, among other benefits. However, not all ridesharing promises have panned out. Recent studies demonstrate that the expected drop in traffic congestion and reduction in greenhouse gas (GHG) emissions have not materialized. This is primarily due to the substantial distances traveled by the ridesharing vehicles without passengers between rides, known as deadhead miles. Recent work has focused on reducing the impact of deadhead miles while considering additional metrics such as rider waiting time, GHG emissions from deadhead miles, or driver earnings. Unfortunately, prior studies consider these environmental and equity-based metrics individually despite them being interrelated.
In this paper, we propose a Learning-based Equity-Aware Decarabonization approach, LEAD, for ridesharing platforms. LEAD targets minimizing emissions while ensuring that the driver’s utility, defined as the difference between the trip distance and the deadhead miles, is fairly distributed. LEAD uses reinforcement learning to match riders to drivers based on the expected future utility of drivers and the expected carbon emissions of the platform without increasing the rider waiting times. Extensive experiments based on a real-world ride-sharing dataset show that LEAD improves fairness by 2$\times$ when compared to emission-aware ride-assignment and reduces emissions by 70% while ensuring fairness within 66% of the fair baseline. It also reduces the rider wait time, by at least 40%, compared to various baselines. Additionally, LEAD corrects the imbalance in previous emission-aware ride assignment algorithms that overassigned rides to low-emission vehicles.
1 Introduction
Road transportation significantly contributes to global energy consumption and greenhouse gas (GHG) emissions [8]. Worldwide, the transportation sector is the fourth-largest source of GHG emissions and is the largest in the United States [42, 29]. In 2021, transportation accounted for 37% of global CO2 emissions [17]. That same year, the United States emitted approximately 5 billion metric tons of carbon dioxide, accounting for about 13.49% of global emissions, surpassing the combined emissions of all 28 European Union countries [7]. Due to these environmental impacts, there is an effort towards exploring transportation systems that transport people in an eco-friendly manner [33].
In recent years, shared mobility options have emerged as an attractive transport solution, thanks to the widespread integration of smartphones into daily life. These services enable users to access transportation on demand without owning a vehicle [34, 4, 26, 27]. As a result, ridesharing is seen as a sustainable transportation option that can help reduce emissions and traffic congestion. Unfortunately, these expectations have not materialized and ridesharing services have been linked to increased traffic congestion, higher emissions, distracted driving, and negative social equity implications [14, 30, 3]. The increase in traffic congestion and emissions results from two key by-products of ridesharing. First, ridesharing availability can reduce the use of public transportation. Second, ridesharing vehicles travel a significant fraction of their total miles without a passenger on their way to pick up a passenger or come back after dropping one [23, 36, 31]. These non-passenger miles are referred to as deadhead miles. Previous studies show that deadhead miles can account for 19% to 41% of total miles while increasing emissions by up to 90% as compared to the personal vehicle use [15, 46].
There has been significant recent work on optimizing the deadhead miles in ridesharing ecosystems [18, 32] while considering additional rider-specific metrics such as the wait time [47, 19] or platform-wide objectives such as GHG emissions [32, 49]. Other studies aim to enhance overall rider satisfaction by minimizing travel costs and waiting times [6, 20, 37]. However, these studies often ignore fairness from the drivers’ perspective, which can exacerbate social inequities. The studies from a diver’s perspective typically focus on enhancing driver satisfaction and operational efficiency by optimizing routes and maximizing a given definition of utility [21, 40]. While some of those studies also target fairness in earnings across drivers [35], they ignore the rider’s wait time and system-wide emissions. As a result, the ride-assignment algorithms proposed by prior cannot be used in practice to promote sustainable and equitable growth in ridesharing platforms.
In this work, we consider the problem of holistic ride assignment that balances the objectives of the system (low emissions), the driver (fairness in earnings), and the rider (low waiting time). The goal is to minimize system-wide emissions, including emissions from ride and deadhead miles, while maximizing fairness in total utility across all trips between drivers without impacting the rider’s waiting time. The utility of a trip for a driver is defined as the trip distance minus the deadhead distance. Given this setup, [32] and [35] represent our most relevant related works. In [32], the authors propose an online threshold-based ride assignment algorithm (TORA) that targets minimizing the system-wide emissions while optimizing the waiting time of the riders. TORA achieves a highly favorable tradeoff between emissions (up to 60% reduction) and rider’s wait time (4% increase). However, to reduce emissions, TORA assigns a disproportionate fraction of rides to low-emission vehicles, resulting in high unfairness among drivers. For example, it may assign up to 65% of the rides to EVs despite them constituting only 25% of the fleet. In [35], the authors propose LAF that targets maximizing utility, defined as the total revenue generated by the system while minimizing unfairness in the driver’s equitable earnings. While taking a driver-centric approach, LAF completely ignores system-wide emissions and the rider’s wait time.
In this paper, we propose the Learning-Based Equity-Aware Decarbonization (LEAD) algorithm, which aims to minimize total system-wide emissions and maximize fairness in drivers’ accumulated utilities across trips. To reduce system-wide emissions, LEAD must consider not only the drivers that are available at present but also the ones that will become available in the near future. To do so, LEAD explicitly accounts for the dependencies among ride assignments by using reinforcement learning to develop future-aware ride assignment strategies. Using a real-world ridesharing dataset, we analyze how LEAD and other representative baselines [32, 35] perform using metrics, such as the reduction in emissions, the quality of service for riders measured as wait time, and fairness for drivers, and how gracefully they navigate the trade-offs between these metrics. Our contributions can be summarized as follows:
  • We frame the problem of ride assignment in ridesharing services (termed RARS) as a bipartite matching problem and propose a new objective function that aims to minimize total carbon emissions across all trips while reducing the gap between utilities of different drivers.
  • We present LEAD, an online learning-based algorithm that reduces emissions and ensures fairness for drivers, as measured by utility. To the best of our knowledge, this is the first study to leverage the dependencies between past and future ride assignments to minimize carbon emissions while ensuring a fair distribution of utility among drivers.
  • We implemented LEAD on a simulation testbed and evaluated its performance using real data from RideAustin [28], a nonprofit ridesharing service. In addition, we compare the performance of LEAD with heuristic and state-of-the-art algorithms used in [32, 35]. Our experimental results show that, in practical scenarios, our algorithm not only significantly improves the fairness of utility distribution among drivers but also substantially reduces emissions, compared to the current state-of-the-art methods.
2 Related Work
Ridesharing Optimization Approaches. The development of ridesharing systems has introduced unique challenges, with the primary focus being the assignment of ride requests to drivers. Some challenges in this area include improving vehicle utilization across multiple ride requests and demands [48, 25, 22], developing algorithms for route planning [10, 43], and analyzing and integrating temporal and spatial patterns to predict the arrival time of riders’ requests [16, 11]. Some other works in this domain also consider different optimization objectives such as maximizing drivers’ profit [45, 2, 51], while others aim to minimize travel costs [41, 44] or reduce riders’ waiting times [50, 1, 47].
In [18] , authors attempt to reduce trip-level deadhead miles by leveraging hour-ahead trip demand predictions and using a heuristic approach to driver assignment. However, this focus on reducing trip-level deadhead miles does not always result in a reduction in system-wide deadhead miles and emissions, as these outcomes also depend on factors such as the fuel efficiency of vehicles and traffic conditions. Another work  [32] focuses on minimizing emissions from deadhead miles while also reducing rider waiting times, ensuring no degradation in user satisfaction. However, a ride assignment system focused only on reducing emissions can negatively impact the quality of service (QoS) for riders and drivers. For instance, prioritizing emission reduction as the primary objective can lead to a preference for electric or low-emission vehicles over high-emission ones. This bias is unfair to drivers with high-emission vehicles, who often belong to lower-income communities and may face reduced earning opportunities. Additionally, an emissions-aware ride assignment may allocate trips with inherently longer deadhead miles to low-emission vehicles. This increases the deadhead-to-trip ratio, decreasing these vehicles’ overall efficiency and service quality.
Another key challenge is the lack of consideration for the temporal dependencies between current and future assignments. This oversight can impact the optimization of utility and fairness. Many algorithms in this area attempt to provide theoretical performance guarantees, but they typically rely on myopic assumptions [47]. For instance, they frequently neglect the dependency of assignments on past and future decisions, which can lead to sub-optimal assignments. This shortsightedness fails to account for the dynamic and evolving nature of ridesharing systems, where current ride assignments can significantly impact future states and the system’s overall efficiency. An effective alternative approach involves leveraging historical data to optimize ride assignments alongside integrating reasonable assumptions.
Fairness in Ridesharing. Fairness is a critical factor in matching scenarios, particularly in multi-sided markets where preferences play a significant role [12, 24, 9]. In the context of ridesharing, recent literature has raised concerns about the fairness of the assignments. Brown et al. [5] shows the unfair treatment of riders by ridesharing companies, which leads to higher trip cancellation rates for a few groups of riders. Another study indicates that income inequality among ridesharing drivers may prevent them from earning a living wage [13]. Another work aims to optimize long-term efficiency and fairness in ridesharing platforms through joint order dispatching and driver repositioning [38]. Another recent work [32] showed a trade-off between reducing carbon emissions and sacrificing the fairness of utilities among drivers. Similarly, another study [35] demonstrated optimization in driver utilities and fairness without considering emissions. This study used reinforcement learning techniques to evaluate the expected achievable utilities in different city areas. Despite the success of this approach in ensuring equity among drivers’ utilities, it did not address eco-friendly ride assignments with low carbon emissions.
Ensuring fair assignments in ridesharing systems is challenging due to drivers’ and riders’ inherent spatiotemporal dynamics. Drivers can earn varying amounts of utility based on their assignments, while riders may experience differing waiting times. These disparities can lead to perceptions of unfairness, complicating the assignment process. To tackle these challenges, our work introduces a new fairness metric for driver earnings that considers the spatiotemporal dynamics inherent in ridesharing systems. This metric is designed to ensure a more equitable distribution of earnings among drivers, considering the dynamic nature of the system. Building on this metric, we have developed a reinforcement learning-based approach assignment algorithm that reduces carbon emissions while providing a fair distribution of driver utilities compared to the state-of-the-art assignment algorithm [32]. The algorithm leverages historical data and dynamic learning to create effective and equitable assignment strategies, addressing immediate and long-term objectives in the ridesharing ecosystem.
3 Problem Formulation
We consider a ridesharing system consisting of a set of drivers, denoted as $\mathcal{V}$, and a set of ride requests, denoted as $\mathcal{R}$. In the ridesharing context, a ride $r\in\mathcal{R}$ is defined as a request that includes a rider’s pickup location, denoted as $p_{r}$, a rider’s dropoff location, denoted as $q_{r}$, and time of the request, denotes as $c_{r}$. A driver $v\in\mathcal{V}$ in the ridesharing system is defined by their current location, utility, and the emissions per mile of the driver’s vehicle $e_{v}$.
In ridesharing platforms, riders set the pickup and dropoff locations for a ride request; then, the platform matches the request to one of the available drivers and estimates the trip duration and waiting time based on factors such as trip distance, traffic congestion, and the distance between the assigned driver and the rider’s pickup location. An available driver in the system refers to a driver who is either idle or soon to be available within the specified time to serve a request. The emission associated with a ride request $r$ is influenced by the unit emissions of the driver’s vehicle, the deadhead distance, the trip distance, and additional factors such as traffic congestion, road conditions, etc. For simplicity, we estimate the carbon emissions produced during the servicing of the ride request $r$ by driver $v$ using the equation below.
$E_{v,r}=(d_{v,r}^{(D)}+d_{r}^{(T)})\cdot e_{v},$ (1)
where $e_{v}$ is the carbon emission per mile of the vehicle of driver $v$, $d_{v,r}^{(D)}$ is the deadhead distance of the trip, and $d_{r}^{(T)}$ is the trip distance which is the distance between rider’s pickup and dropoff location. The utility of a driver $v$ is defined as the difference between the trip distance and deadhead distance accumulated across all served ride requests. The trip distance represents the actual distance driven to fulfill the ride request, which contributes to the driver’s earnings. In contrast, the deadhead distance, the distance traveled without carrying a passenger (e.g., traveling from the driver’s location to the rider’s pickup location), incurs an overhead cost for the driver beyond the costs incurred during the trip, such as fuel expenses. By subtracting the deadhead distance from the trip distance, we effectively account for the revenue generated and the overhead cost incurred by the driver. This approach provides a measure of the driver’s net revenue, reflecting the true utility of a ride to the driver:
$U_{v,r}=\sum_{r^{\prime}\in\mathcal{V}_{r}}(d_{r^{\prime}}^{(T)}-d_{v,r^{\prime}}^{(D)})\cdot x_{v,r^{\prime}},$ (2)
where $U_{v,r}$ is the cumulative utility of the driver $v$ before serving ride request $r$, $\mathcal{V}_{r}$ is the set of ride requests posted before $r$, and $x_{v,r}\in\{0,1\}$ is a decision variable: $x_{v,r}=1$ if ride request $r$ is assigned to driver $v$; 00 otherwise. For simplicity, we denote the utility of driver $v$ after serving all requests as $U_{v}$.
Table 1: Summary of important notations.
Notation Description
$\mathcal{V}$ Set of all drivers
$\mathcal{R}$ Set of all ride requests
$\mathcal{V}_{b}$ Set of available drivers during batch $b$
$\mathcal{R}_{b}$ Set of ride requests in batch $b$
$p_{r}$ Pickup location of ride request $r$
$q_{r}$ Dropoff location of ride request $r$
$c_{r}$ Created time of ride request $r$
$a_{r,v}$ Dropoff time of ride request $r$ when served by driver $v$
$E_{v,r}$ Emission of serving request $r$ by driver $v$
$U_{v,r}$ Utility history of driver $v$ before serving ride request $r$
$d^{(T)}_{r}$ Trip distance of ride request $r$
$d^{(D)}_{v,r}$ Deadhead distance of serving ride request $r$ by driver $v$
$d_{t,D}$ Deadhead distance of $t^{th}$ trip served by the agent
$d_{t,T}$ Trip distance of $t^{th}$ trip served by the agent
$V_{D}(s)$ Value function at state $s$ associated to deadhead distances
$V_{T}(s)$ Value function at state $s$ associated to trip distances
$e_{v}$ Produced carbon emission per mile of the vehicle of driver $v$
The ride-assignment problem builds on the classic bipartite matching problem, involving riders and drivers as distinct sets. The main challenge is modeling the constraints of available drivers and their costs for new requests, especially when some drivers are already occupied but may soon become available to take a ride. Our goal is to design an algorithm that assigns available vehicles to ride requests in a manner that minimizes emissions while ensuring fairness among drivers and considering the long-term effects and interdependencies of current assignments on future ones. We formulate the RARS problem as follows.
[…]
$\text{Emission}:=\sum_{v}\sum_{r}E_{v,r}\cdot x_{v,r},$
$\text{Fairness}:=-\max_{v}(U_{v})$
$\quad\ \ \ \ +\min_{v}(U_{v}),$
s.t., $\sum_{v}x_{v,r}\in\{0,1\}\qquad\forall r,$ (3e)
$\sum_{\mathclap{\{r^{\prime}\neq r|c_{r}\leq c_{r^{\prime}}\leq a_{r,v}\}}}x_{v,r^{\prime}}\cdot x_{v,r}=0\quad\ \forall v,r,$
vars., $\ \ \quad x_{v,r}\in\{0,1\}\qquad\forall v,r.$ (3f)
where $a_{r,v}$ shows the dropoff time of ride request $r$ when served by driver $v$. The term Emission includes trip emissions and deadhead miles emissions. The Fairness metric is defined as the difference between the minimum and maximum utilities of the drivers and is a non-positive metric. Constraint (3e) ensures that each ride request is assigned to at most one driver, and constraint (3e) ensures that each driver is assigned to at most one ride request at each time. The parameter $\eta\geq 0$ controls the balance between emissions and fairness. Specifically, it quantifies the increase in emissions in grams of carbon dioxide (g.CO2) that we will accept for a one-kilometer decrease in unfairness across drivers. A higher value of $\eta$ indicates a greater emphasis on fairness. A lower value of $\eta$ prioritizes reducing emissions, potentially at the cost of sacrificing fairness.
4 The LEAD Algorithm
This section introduces the LEAD algorithm that leverages reinforcement learning and batching to accumulate multiple ride requests that are jointly assigned to ensure that a long-term system-wide beneficial assignment is obtained. We present the formulation of the single batch matching problem as an integer linear programming problem whose objective includes two terms: one for the expected emission of ridesharing service and the second for the expected fairness in utility across drivers. To evaluate these expected values, LEAD employs reinforcement learning to account for temporal dependencies between current and future ride-matching. This is essential since the system operates online, with rider requests and driver availability dynamically changing in real-time.
In the following, we present the details of LEAD, which consists of two main modules: the Learning Based module and the Batched Emission Aware Fair Assignment module. These modules are discussed in detail in Section 4.1 and Section 4.2, respectively.
Figure 1: Overview of LEAD. During each batch, LEAD evaluates the long-term emissions and utilities and uses them to construct the weighting of the ILP problem of (10). After solving the optimization problem and matching ride requests with drivers, LEAD updates the value functions based on the deadhead and trip distances of the served requests in the batch.
4.1 Learning Expected Emissions and Utility
In the following, we discuss the use of online reinforcement learning to model the impact of current assignments on future assignments and evaluate the expected deadhead and trip distances of trips in different regions of the covered area. Reinforcement learning is a method where agents learn by interacting with the environment over time. At each step, the agent takes an action and receives feedback in the form of utility from the environment. This continuous interaction allows the agent to optimize its strategy based on the utility received, making it well-suited for dynamic and complex systems such as ridesharing, where decisions made at one point can influence outcomes in the future. In our work, the reinforcement learning module consists of the following elements:
  • Agent: Each active driver $v$ in the ridesharing system is considered as an agent.
  • State: The state of driver $v$ during batch $b$ is represented by the driver’s location $l_{v,b}$. To facilitate data analysis and management, city space is typically discretized into a square grid system, which results in a finite number of unique states.
  • Action: The action of the available driver is defined as the request they choose to accept.
After taking action, each agent receives utility, which is the difference between the trip distance and the deadhead distance. To predict the expected emissions produced and the utility achieved by a driver, we introduce two value functions for the agent: one for the deadhead distance and another for the trip distance.
Due to the complexity of considering multiple dependent agents within the system, we simplify the learning module by treating different agents as a single agent. In the emission-aware fair assignment module, we leverage the evaluated value functions to take appropriate actions for the different action-correlated agents. We adopt a framework that involves sequences comprising the current state, current action, reward, and next state to estimate the value functions. One effective method for learning in this scenario is Temporal Difference (TD) learning [39], which iteratively updates value functions based on the deadhead and trip distances of each completed ride request.
Below, $V_{D}(s)$ shows the expected deadhead distance a driver could travel when starting from state $s$.
$V_{D}(s)=\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}\cdot d_{t,D}|s_{0}=s\right],$ (4)
where $d_{t,D}$ is the deadhead distance of $t^{th}$ trip served by the agent, $s_{0}$ denotes the starting state of the agenet, and $\gamma\leq 1$ is the discount factor. Similarly, we define $V_{T}(s)$ as the expected trip distance a driver will travel when starting from state $s$.
$V_{T}(s)=\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}\cdot d_{t,T}|s_{0}=s\right],$ (5)
where $d_{t,T}$ is the trip distance of $t^{th}$ trip served by the agent.
Assume the agent is located in state $s_{t}$ and after serving the request, travels to the next state $s_{t+1}$. The TD learning update rule for the value functions based on the trip and deadhead distance of the trip are as follows:
$V_{D}(s)\leftarrow$ $V_{D}(s)+\alpha\left[d_{t,D}+\gamma V_{D}(s_{t+1})-V_{D}(s_{t})\right],$ (6)
$V_{T}(s)\leftarrow$ $V_{T}(s)+\alpha\left[d_{t,T}+\gamma V_{T}(s_{t+1})-V_{T}(s_{t})\right],$ (7)
where $\alpha$ is the learning rate. The values of $V_{D}(s)$ and $V_{T}(s)$ are initialized through a specific initialization process (such as zero initialization or random initialization). After a driver serves a ride request, the values of these functions are updated accordingly. In the next section, we demonstrate how these value functions can be used to evaluate the expected deadhead and trip distances for trips starting in different city locations, ultimately estimating the future emissions and utility of drivers.
1 forall $v\in\mathcal{V}_{b},r\in\mathcal{R}_{b}$ do
2       Calculate $E_{v,r,b}$ using Equation (4.2);
3       Calculate $U_{v,r,b}$ using Equation (4.2);
4
5Matched-Pairs $\leftarrow$ Solution of optimization problem (10);
6 foreach $r\in\mathcal{R}_{b}$ do
7       $v$leftarrow Matched-Pairs($r$);
8       $s_{t}\leftarrow$ current location of $v$;
9       $s_{t+1}\leftarrow$ dropoff location of the request $r$;
10       Update $V_{D}(s_{t})$ using Equation (6);
11       Update $V_{T}(s_{t})$ using Equation (7);
12
13return Matched-Pairs;
Algorithm 1 LEAD($\mathcal{R}_{b}$, $\mathcal{V}_{b}$)
4.2 Batched Emission-Aware Fair Assignment
In this section, we propose the process of matching ride requests to available drivers within LEAD. LEAD performs the ride matching process in a batch setting, where each batch comprises ride requests posted after the assignments from the previous batch. Let $E_{v,r,b}$ denote the long-term expected emission produced by driver $v$ after being assigned to request $r$ within batch $b$. The value of $E_{v,r,b}$ depends on the trip and deadhead distances for serving request $r$, as well as future deadhead and trip distances traveled by driver $v$. To estimate the latter, we leverage the value function described in the previous section. Specifically, the expected emission produced by driver $v$ during and after batch $b$ can be calculated as:
$E_{v,r,b}$ $=\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}(d_{t,T}+d_{t,D})\cdot e_{v}\big{|}s_{0}=l_{v,b},s_{1}=q_{r}\right]$
$=\bigg{[}(d_{r}^{(T)}+d_{v,r}^{(D)})+\gamma(V_{T}(q_{r})+V_{D}(q_{r}))$
$-(V_{T}(l_{v,b})+V_{D}(l_{v,b}))\bigg{]}\cdot e_{v},$ (8)
which is directly derived from Equations (4) and (5), considering that the driver is located at state $l_{v,b}$ and will travel to the dropoff location of the request after serving of that. Similarly, let $U_{v,r,b}$ denote the expected final utility of driver $v$ given that they were assigned to request $r$ during batch $b$ and had a utility history of $U_{v,r}$ by that time. We can evaluate the value of $U_{v,r,b}$ as follows:
$U_{v,r,b}$ $=\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}(d_{t,T}-d_{t,D})\big{|}s_{0}=l_{v,b},s_{1}=q_{r}\right]$
$=U_{v,r}+(d_{r}^{(T)}-d_{v,r}^{(D)})+\gamma(V_{T}(q_{r})-V_{D}(q_{r}))$
$-(V_{T}(l_{v,b})+V_{D}(l_{v,b})).$ (9)
During the ride-matching of batch $b$, LEAD finds the set of requests $\mathcal{R}_{b}$, and available drivers in that batch, $\mathcal{V}_{b}$, and evaluates $E_{v,r,b}$, and $U_{v,r,b}$ for every pair of $v$ and $r$ in the batch. Then, LEAD finds the solution of the integer linear programming problem below to match each ride request to available drivers.
$\min\quad$ $\text{Emission}_{(b)}-\eta\cdot\text{Fairness}_{(b)},$ (10a)
$\text{Emission}_{(b)}:=\sum_{v\in\mathcal{V}_{b}}\sum_{r\in\mathcal{R}_{b}}E_{v,r,b}\cdot x_{v,r},$ (10b)
$\text{Fairness}_{(b)}:=-\max_{v}\sum_{r\in\mathcal{R}_{b}}(U_{v,r,b}\cdot x_{v,r})$
$+\min_{v}\sum_{r\in\mathcal{R}_{b}}(U_{v,r,b}\cdot x_{v,r}),$ (10c)
s.t., $\sum_{v\in\mathcal{V}_{b}}x_{v,r}\in\{0,1\}\quad\forall r,$ (10d)
$\sum_{r\in\mathcal{R}_{b}}x_{v,r}\in\{0,1\}\quad\forall v,$ (10e)
vars., $\quad\quad\ x_{v,r}\in\{0,1\}\quad\forall v,r,$ (10f)
where $\text{Emission}_{(b)}$ represents the expected total emissions produced by the ridesharing service during and after batch $b$. $\text{Fairness}_{(b)}$ estimates the expected fairness of the system given the assignments in batch $b$. It is worth mentioning that if the number of ride requests in batch $b$ exceeds the number of available drivers, only the first $|\mathcal{V}_{b}|$ ride requests will be assigned, with the remaining requests being moved to the next batch.
The pseudocode for LEAD is outlined in Algorithm 1. Initially, LEAD evaluates the values of $E_{v,r,b}$ and $U_{v,r,b}$ for every pair of $v$ and $r$ in batch $b$ (Lines 1-3). Next, it finds the solution for the optimization problem in (10), determining which driver should serve each ride request. The final step is to update the value functions based on the actual trip and deadhead distances of the served ride requests within this batch. Specifically, for each request $r$ in the batch, LEAD updates the value functions associated with trip and deadhead distances using the update rules presented in Equations (6) and (7) (Lines 5-10). By integrating emission and fairness into a single objective, LEAD balances reducing carbon emissions with the fairness of utilities among drivers. Additionally, by considering the long-term expected emissions and utilities instead of just the current values, LEAD optimizes these metrics with a forward-looking perspective.
5 Experimental Evaluation
In this section, we conduct comprehensive experiments to evaluate the performance of LEAD using several metrics, including emissions per trip, fairness in driver’s utility, and average wait times for riders. We compare LEAD against state-of-the-art algorithms across various values for different batch durations, the fairness parameter $\eta$, and the percentages of low-emission vehicles. Below, we outline the key questions the evaluation addresses and summarize our findings.
  1. Q1
    How does LEAD reduce emissions, especially compared to other baselines targeting emissions reduction?
    Outcome: LEAD reduces emissions by at least 73.6%, 70.5%, and 58.3% compared to CD, TORA, and LAF algorithms, respectively.
  2. Q2
    How much unfairness does LEAD incur to reduce emissions, compared to LAF that only targets fairness?
    Outcome: LEAD offers a favorable, and more importantly configurable, tradeoff between fairness and emissions reduction; it reduces emissions by 65% for a 9% drop in fairness. It also exposes the $\eta$ knob to enable navigating the tradeoff gracefully.
  3. Q3
    How does LEAD’s rider wait time performance compare to the baselines that explicitly consider reducing wait time?
    Outcome: LEAD achieves at least a 40% improvement over baseline algorithms while also ensuring that improving fairness for drivers does not impact the rider’s wait time by more than 3%.
Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
(a) LEAD vs. baselines (b) Impact of LEVs (c) Impact of Eta ($\eta$).
Figure 2: Emissions reduction performance: (a) emission per trip for different algorithms as a function of batch duration, (b) impact of the percentage of low emission vehicles (LEVs) in the fleet on LEAD, and (c) the impact of increasing emissions for an increase in fairness, captured using $\eta$. Here, $\eta$ specifies extra emissions that the algorithm incurs to reduce unfairness by 1km.
5.1 Experimental Setup
First, we provide details on the ridesharing dataset, baseline algorithms, metrics, and the range of evaluation parameters.
Ridesharing dataset. Our experiments use a publicly available dataset from RideAustin [28], a non-profit ridesharing service based in Austin, Texas. This dataset includes approximately 1.5 million trips over ten months in 2016 and 2017, collected in Austin. It contains detailed trip information such as pick-up and drop-off coordinates of each trip, vehicle make and model, distances traveled before, during, and after each trip.
In our experiments, we use a subset of the RideAustin dataset, focusing on trips from December 1, 2016, to December 15, 2016. This subset includes 58,866 ridesharing trips served by 150 drivers. We augment this dataset by including carbon emission data for both deadhead miles and individual trips, as well as equity information for drivers and riders, using E2-RideKit, a publicly available toolkit [32]. We assume a rider cancels their request if it remains unassigned for two consecutive batches, and considered drivers available within 15 minutes in the set of available drivers of the batch. We also classified vehicles emitting under 135 g.CO2/km as low-emission (LEVs) and over 270 g.CO2/km as high-emission (HEVs). To assess the impact of replacing HEVs with LEVs, we created three dataset variants by randomly converting a certain percentage of the non-LEVs into electric vehicles (EVs), resulting in datasets with 10%, 15%, and 25% LEVs. We set the EV’s emission intensity to 63.35 g.CO2 per kilometer. [^1] We used the Tesla Model Y as a representative EV, which has an energy efficiency rating of 26 kWh/100mi. The unit distance emissions were calculated using the average carbon intensity of 408 g.CO2eq/kWh for Austin, Texas.
Comparison algorithms. We evaluate the performance of LEAD against a heuristic baseline algorithm and two state-of-the-art algorithms from prior works detailed below:
  • Closest Driver (CD): For a given batch of requests, this algorithm sequentially matches requests in the batch and greedily assigns a request to the nearest available driver to minimize the deadhead miles and the rider’s wait time.
  • LAF [35]: This algorithm has four stages: evaluating, assigning, guiding, and learning. Initially, it evaluates a bipartite graph where edge weights represent trip prices, and these weights are updated to reflect both current and future earnings. During the assigning stage, LAF utilizes a bi-objective assignment algorithm to balance the aggregate utilities of drivers and the equity in their utilities. The learning stage involves refining weight updates through reinforcement learning to guide future assignments.
    Since LAF aims to maximize total utilities and ensure equity in utilities, it uses a single value function to track the expected achievable utilities in different city regions. For adaptation to the current work, we set the utility (or price) of a ride request $r$ served by driver $v$ as the difference between the trip and deadhead distances and evaluated weights in Algorithm 1 of [35] accordingly. Additionally, during our experiments, the weight of trip utilities over different times of the day is the same.
  • TORA  [32]: TORA is an online ride-assignment algorithm that aims to balance passenger waiting times with system-wide emissions. To the best of our knowledge, it is the only existing online ride assignment algorithm that explicitly considers emissions. TORA aims to reduce passenger waiting times by initially selecting the closest available driver. It then evaluates alternative available drivers by comparing their distances and the additional deadhead emissions they would generate relative to the closest driver. For each driver, TORA calculates an “Emission-to-Distance” (E2D) ratio, which is the ratio of the difference in deadhead emissions and the deadhead miles from the two drivers to the passenger. Finally, the algorithm selects the driver with the most favorable E2D ratio. Although TORA accounts for carbon emissions during ride matching, it does not address fairness in driver utilities.
Parameter ranges. We evaluate the performance of LEAD and baseline algorithms in various experimental settings. We evaluate the impact of parameter $\eta$ and batch duration by varying $\eta$ between 0.1 g.CO2/km and 10 g.CO2/km and varying batch duration values between 2 and 15 minutes. To define unique states, we divide the city into square tiles with a width of 1 km, initialize value functions to 0, set the discount factor $\gamma=0.9$, and the learning rate $\alpha=0.025$ to align with the evaluation settings used in [35].
Performance metrics. We evaluate the performance of LEAD and comparison algorithms using three performance metrics: per trip emissions, fairness in driver’s utility, and rider waiting time. For fairness, we normalize our results to LAF outcomes when comparing different algorithms. However, in analyzing the effect of LEV penetration, we normalize the values to the LEAD without the addition of LEVs. We also report the proportion of trips assigned to low-emission vehicles (LEVs) and high-emission vehicles (HEVs), as well as the proportion of successfully matched requests.
Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
(a) LEAD vs. baselines (b) Impact of LEVs (c) Impact of Eta ($\eta$).
Figure 3: Fairness performance: (a) normalized fairness for different algorithms as a function of batch duration, (b) impact of the percentage of low emission vehicles (LEVs) in the fleet on LEAD, and (c) the impact of increasing emissions for an increase in fairness, captured using $\eta$. Here, $\eta$ specifies extra emissions that the algorithm incurs to reduce unfairness by 1km.
5.2 Impact on Per-Trip Emissions
Figure 2 shows the performance of LEAD in reducing the average emissions per trip against the baselines across various parameters.
Key results. As shown in Figure 2(a), LEAD outperforms all the baseline algorithms for all batch durations. LEAD reduces emissions per trip by at least 73.6%, 70.5%, and 58.3% when compared to CD, TORA, and LAF algorithms, respectively. LEAD does even better when the batch duration increases. While the emissions per trip for CD and TORA slightly decrease with an increase in batch duration, LEAD achieves significantly higher reductions at longer batch duration, thereby increasing the margin of improvement. Finally, LAF’s performance worsens with an increase in batch duration; emissions are 9% and 3% higher at 15 minutes than at 2 and 5 minutes, respectively.
As shown in Figure 2(b), adding low-emission vehicles (LEVs) to the ridesharing fleet reduces the average emissions per trip. Adding 10% LEVs provides up to 8.2% decrease in emissions; adding 15% LEVs reduces emissions by up to 13.5%; adding 25% LEVs achieves up to 30.9% reductions in emissions. Importantly, the lowest reductions for all LEV values occur at the longest batch duration, such as adding 10% LEVs at 15 minutes yields no added benefit.
As shown in Figure 2(c), $\eta$ does not have a significant impact on the emissions per trip. The highest increase in the emissions per trip is 5.15%, which is observed for an LEV percentage of 25%. As we increase $\eta$ to 10gCO2/km, the emissions only increase by 1.9% and 2.5% at 10% and 15% LEV percentages, respectively.
Analysis of findings. We next take a deeper look at some of the results we have observed about LEAD’s impact on emissions.
  • Why does LEAD outperforms deadhead miles- and emissions-aware algorithms CD and TORA? The primary reason that LEAD outperforms CD is that CD does not take into account the unit emissions of the vehicles when assigning requests to drivers. Another reason that LEAD outperforms TORA and CD is that it assigns all the requests in each batch simultaneously, whereas TORA and CD assign the requests sequentially within each batch.
  • Why does adding LEVs yield fewer emissions reductions at longer batches? As the batch duration increases, LEAD can access a bigger set of drivers and riders, allowing it to make low-emission ride assignments, even without a high LEV percentage. The effect is more acute at smaller percentages of LEVs.
  • Why do the emissions per trip for LEAD not increase with $\eta$? The value of $\eta$ shows the grams of CO2 emitted to increase fairness by 1km. LEAD makes this tradeoff highly judiciously and only chooses the trips when improvement in fairness is possible.
Key takeaways. LEAD outperforms the state-of-the-art emission-aware ride assignment algorithm TORA by up to 80.7%. While additional reductions may be possible by optimizing batch duration and $\eta$, it will introduce tradeoffs with our other objectives of increasing driver’s fairness and reducing rider’s wait time.
Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
(a) LEAD vs. baselines (b) Impact of LEVs (c) Impact of Eta ($\eta$).
Figure 4: Wait time performance: (a) wait time for different algorithms as a function of batch duration, (b) impact of the percentage of low emission vehicles (LEVs) in the fleet on LEAD performance, and (c) the impact of increasing emissions for an increase in fairness, captured using $\eta$. Here, $\eta$ specifies extra emissions that the algorithm incurs to reduce unfairness by 1km.
5.3 Impact on Driver’s Fairness
Figure 3 shows the performance of LEAD in improving fairness among drivers against the baselines across various parameters.
Key results. As shown in Figure 3(a), no other algorithm, including LEAD, beats the fairness performance of LAF. The next best fairness performance is shown by LEAD, as it reaches 90.8% of the fairness achieved by LAF at the batch duration of 5 minutes. Importantly, LEAD’s worst performance of achieving 65% of LAF’s fairness is only marginally surpassed by CD algorithm which achieves 69% of LAF’s fairness performance. TORA is the worst performing algorithm as it achieves between 37% to 49% of LAF’s fairness performance.
As shown in Figure 3(b), adding LEVs gives the greatest benefits at low batch duration, where having 10% and 15% LEVs can improve fairness by 36% and to almost 2$\times$, respectively, compared to baseline ridesharing fleet with LEAD. Increasing LEVs beyond 15% provides marginal gains in fairness across all batch durations. Finally, the gains in fairness have the lowest point at a medium batch duration of 10 minutes and increase on either side of the batch duration.
Figure 3(c) shows the effect of $\eta$ on fairness, providing us with two key observations. First, increasing $\eta$ increases fairness, with a 37% increase at 0.1g.CO2/km and 80% increase at 10g.CO2/km. Second, even a small percentage of LEVs provide a very big improvement in fairness at the higher values of $\eta$; 15% LEVs can improve fairness by 40% at $\eta$ value of 10g.CO2/km.
Analysis of findings. We next look deeper at the results we presented to quantify the fairness properties of LEAD.
  • Why does fairness first increase and then decrease with increasing batch duration? As the batch duration increases, more opportunities exist to increase fairness and reduce emissions. The absolute fairness (not shown in the graph) for both LEAD and LAF increases. However, since LAF uses the benefits from batch duration only to improve fairness, while LEAD uses it for both increasing fairness and decreasing emissions, the normalized performance of LEAD with respect to LAF decreases.
  • Why does longer batch duration suppress the effect of LEVs on improving fairness? When the batch duration is short, there are fewer available drivers, limiting the algorithm to a smaller pool and reducing its ability to prioritize low-emission vehicles. However, with longer batch durations, the algorithm has a wider selection, allowing it to choose low-emission vehicles more effectively, which can lead to reduced fairness.
  • How increasing $\eta$ to improve fairness impacts emissions? As outlined in the previous section, $\eta$ has a marginal impact on emissions, with a maximum increase of 5.15%. Since increasing $\eta$ can improve fairness by almost 80%, the tradeoff seems desirable.
Key takeaways. LEAD offers a highly favorable tradeoff between improving fairness and reducing emissions. For example, at batch duration of 5 minutes, it reduces emissions by 65% while degrading fairness by less than 9%. However, using $\eta$, fairness can be made comparable with LAF for a less than 2% increase in emissions.
5.4 Impact on Rider’s Wait Time
Figure 4 shows the impact of LEAD on the wait time for riders against the baselines across various parameters. Note that CD targets reduce wait time and deadhead miles (or emission) by assigning the closest driver. TORA is explicitly wait time aware as it bounds the wait time experienced by the rider.
Key results. As shown in Figure 4(a), LEAD outperforms the baseline algorithms for all batch duration. The reduction in wait time over CD is the most significant, with 49% to 54% improvement. The performance improvement over LAF and TORA is always more than 40%. Finally, the increase in batch duration has a higher impact on our baseline algorithms, while LEAD achieves its best wait time performance at a batch duration of 5 minutes.
Figure 4(b) shows that the wait time for the riders gets worse at a higher penetration of LEVs. However, the increase in wait time, even at the worst point, is 4.16% (for 15% LEVs at 5-minute batch duration). Interestingly, increasing LEVs from 15% to 25% reduces the wait time, but by at most 1.6% and at least 0.2%.  Figure 4(c) shows that increasing $\eta$ has a small effect as well, between 2.1% to 2.8% for different penetration levels of LEVs.
Analysis of findings. We now do a deep dive into our results.
  • Why does wait time for LEAD increase after batch duration of 5 minutes? An increase in the batch duration puts a lower bound on the wait time for all the rides in the batch. However, an initial increase in batch size duration increases the possibility of a driver coming closer to the rider and becoming available. However, beyond a certain batch duration, the effect of increasing the lower-bound dominates, and wait time goes up.
  • Why having more LEVs is bad for wait time? How does it impact the usefulness of LEAD? As LEAD targets reducing emissions, it will prioritize EVs even if they are farther away or will only be available after finishing a ride, thus increasing the wait time. While the increase in wait time is not desirable, the increase is not significant, and the addition of LEVs provides significant emission reduction and fairness benefits.
Key takeaways. LEAD outperforms CD by considering the future availability of the drivers, same as TORA. However, its explicit focus on reducing deadhead miles reduces wait time beyond TORA. Importantly, LEAD’ wait time is less sensitive to batch duration and $\eta$, allowing for reducing emissions and increasing fairness.
5.5 Additional Performance Benchmarking
In this section, we benchmark LEAD and other baselines using the additional metrics that evaluate system-level performance. The results are shown in Table 2 and Table 3.
LEAD outperforms baseline algorithms in improving utility by at least 19% and reducing deadhead-to-trip ratio by 15.2% over the second-best algorithm LAF. However, LAF rejects many ride requests, as outlined in the original paper, and accepts only 93.3% requests compared to 98.3% ride acceptance rate for LEAD.
Importantly, LEAD also solves TORA’s problem of assigning a disproportionate number of rides to LEVs by allocating rides to LEVs in proportion to their penetration in the fleet.
Table 2: Performance comparison of LEAD against baselines for $\eta=5$ g.CO2/km and batch duration of $5$ minutes.
Algorithm Avg. utilities (km) Avg. deadhead Percentage of
to trip ratio matched requests
CD 763.14 0.66 $97.4\%$
TORA [32] 820.4 0.62 $97.3\%$
LAF [35] 890.56 0.59 $93.3\%$
LEAD (ours) 1065.09 0.50 $98.3\%$
Table 3: Fraction of rides assigned to LEVs at various LEV %ages for $\eta$=5 g.CO2/km and batch duration of $5$ minutes.
Algorithm Original 10% LEV 15% LEV 25% LEV
TORA [32] 13.58 20.41 26.07 35.05
LEAD (ours) 6.9 11.60 16.61 26.16
6 Discussion and Future Works
In this paper, we propose LEAD, a Learning-based Equity-Aware Decarbonization approach for ridesharing platforms. LEAD minimizes emissions while ensuring fair distribution of drivers utility. LEAD leverages reinforcement learning to optimize rider-driver matches based on the expected future utility for drivers and projected carbon emissions for the platform without increasing rider waiting times. Extensive experiments using a public real-world ridesharing dataset demonstrate that LEAD outperforms state-of-the-art baselines in terms of reduction in emissions and wait time while providing a desirable tradeoff with fairness. In the future, we will explore fairness from a rider’s perspective and short-term fluctuations in driver utility, in addition to the long-term effect this paper considers.
7 Acknowledgements
This research was supported in part by NSF grants CAREER 2045641, CNS-2102963, CNS-2106299, CPS-2136199, NRT-2021693, and NGSDI-2105494.